OpenStudy (anonymous):

d=2t^2-7 how would I change the equation t in terms of d

OpenStudy (whpalmer4):

\[d = 2t^2 - 7\] If you want t in terms of d, add 7 to both sides, divide both sides by 2, then take the square root of both sides. That should give you t in terms of d.

OpenStudy (whpalmer4):

You're just methodically moving everything else to the other side.

OpenStudy (anonymous):

ok yea I did that but I have no idea what to do with the squareroot

OpenStudy (whpalmer4):

\[d = 2t^2 -7\rightarrow d+7 = 2t^2 \rightarrow \frac{d+7}{2} = t^2 \rightarrow \sqrt{\frac{d+7}{2}} = t\]

OpenStudy (whpalmer4):

And because there's a square root, t could = either the positive or negative value from a mathematical standpoint. From a physical standpoint, if this equation represents physical quantities, a negative value for t may not make sense.

OpenStudy (whpalmer4):

Do you see why there can be two different values of t for one value of d? Let's say t = 1, then\[ d = 2(1)^2 - 7 = -5\] But what if t = -1? \[ d = 2(-1)^2 - 7 = -5\]Two identical values of d produced by different values of t.